Given two complex numbers z1=1+i and z2= √(3)-1

a) Write z1/z2 in algebraic and polar forms.

b) Deduce the exact values of Cos 5pie/12 and Sin 5pie/12

c) What is the lowest positive value of integer n such that (z1/z2) is real


1
Expert's answer
2021-11-08T17:39:55-0500

a)


"\dfrac{z_1}{z_2}=\dfrac{1+i}{\sqrt{3}-i}=\dfrac{(1+i)(\sqrt{3}+i)}{3-1}"

"=\dfrac{\sqrt{3}+i+i\sqrt{3}-1}{4}=\dfrac{\sqrt{3}-1}{4}+\dfrac{\sqrt{3}+1}{4}i"

"z_1=1+i, |z|=\sqrt{1^2+1^2}=\sqrt{2}"

"\tan\theta_1=\dfrac{1}{1}=1"

"z_1=1+i=\sqrt{2}(\cos(\pi/4)+i\sin(\pi/4))"

"z_2=\sqrt{3}-i, |z_2|=\sqrt{(\sqrt{3})^2+(-1)^2}=2"

"\tan\theta_2=\dfrac{-1}{\sqrt{3}}, \theta_2=-\pi/6"

"z_2=\sqrt{3}-i=2(\cos(-\pi/6)+i\sin(-\pi/6))"

"\dfrac{z_1}{z_2}=\dfrac{1+i}{\sqrt{3}-i}=\dfrac{\sqrt{2}(\cos(\pi/4)+i\sin(\pi/4))}{2(\cos(-\pi/6)+i\sin(-\pi/6))}"

"=\dfrac{\sqrt{2}}{2}(\cos(\pi/4+\pi/6)+i\sin(\pi/4+\pi/6))"

"=\dfrac{\sqrt{2}}{2}(\cos(5\pi/12)+i\sin(5\pi/12))"


b)


"\dfrac{\sqrt{3}-1}{4}=\dfrac{\sqrt{2}}{2}\cos(5\pi/12)"

"\cos(5\pi/12)=\dfrac{\sqrt{6}-\sqrt{2}}{4}"


"\dfrac{\sqrt{3}+1}{4}=\dfrac{\sqrt{2}}{2}\sin(5\pi/12)"

"\sin(5\pi/12)=\dfrac{\sqrt{6}+\sqrt{2}}{4}"

c)


"(\dfrac{z_1}{z_2})^n=(\dfrac{\sqrt{2}}{2})^n(\cos(\dfrac{5\pi n}{12})+i\sin(\dfrac{5\pi n}{12}))"

If "(\dfrac{z_1}{z_2})^n" is real, then "\sin(\dfrac{5\pi n}{2})=0."

"\dfrac{5\pi n}{12}=\pi m, m\in \Z"

"n=\dfrac{12m}{5}"

The lowest positive value of integer "n" is "n=12."